I need to solve the equations with the smallest denominator.
9/10 - 11/18 =
So, I'm thinking the common denominator is 180 and the answer would be 7/10 but I'm not sure because don't both these numbers have a LCM of 2? If it was 2 I'm not sure how to work it?
I appreciate your help cuz I'm so confused!:)
Thanks,
Skyler

Answer by Alan3354(69443)

I need to solve the equations with the smallest denominator.
9/10 - 11/18 =
So, I'm thinking the common denominator is 180 and the answer would be 7/10 but I'm not sure because don't both these numbers have a LCM of 2? If it was 2 I'm not sure how to work it?
================
1st, it's not an equation, there's nothing on the right side.
2nd, there's nothing to solve for, no variable.
It's an addition problem. (Subtraction is a form of addition.)
9/10 - 11/18
180 is a common denominator
= 162/180 - 110/180
= 52/180
= 13/45
=================
But, you mentioned the "smallest denominator."
Teachers are obsessed with using the least common denominator (LCD) for no good reason.
------
The LCD is 90, not 180
9/10 - 11/18
Using the LCD:
= 81/90 - 55/90
= 26/90
= 13/45
Same answer, but you've wasted time fretting about the LCD. That could make a difference on a timed test.